OpenStudy (anonymous):

A cattle train left Miami and traveled toward New York. 14 hours later a diesel train left traveling at 45 km/h in an effort to catch up to the cattle train. After traveling for four hours the diesel train finally caught up. What was the cattle train's average speed

OpenStudy (anonymous):

Speed = Distance / Time (Distance is the same for both trains) Speed * Time = Distance S1 = s T2 = 14+4 hours = 18 hours S2 = 45 T2 = 4 s*18 = 45*4 s = 180/18 s = 10km/h